# Graph sensitivity under join and decomposition

### v2

@@ 1. Introduction
The idea of graph sensitivity traces back to work related to Boolean functions and the hypercube graph ([Gotsman-Linial-92, CFGS-88]), but has been formalized and studied for other graphs following Huang’s clever and concise resolution in [Huang-19] of the sensitivity conjecture for Boolean functions [Nisan-Szegedy-94]. The hypercube graph, $Q_{n}$, which has vertices the $n$-tuples in $\mathbb{Z}_{2}^{n}$ with two vertices adjacent if they differ in exactly one coordinate, is bipartite so its independence number is $\alpha(Q_{n})=\frac{1}{2}2^{n}=2^{n-1}$. Huang proved the smallest maximum degree of an induced subgraph of $Q_{n}$ on $\alpha(Q_{n})+1$ vertices is at least $\sqrt{n}$ and the construction in [CFGS-88] implies it is equal to $\left\lceil\sqrt{n}\,\right\rceil$. Huang further suggests studying, for other highly symmetric graphs, the smallest maximum degree of an induced subgraph of a graph $G$ on $\alpha(G)+1$ vertices, defined as the sensitivity $\sigma(G)$ in [GarciaMarco-Knauer-22] by Garcìa-Marco and Knauer.
@@
Since $Q_{n}$ is the Cayley graph on $\mathbb{Z}_{2}^{n}$ with connection set $S=\{e_{1},\dots,e_{n}\}$ it is natural to investigate the sensitivity of other Cayley graphs. For example, in [Alon-Zheng-20] Alon and Zheng proved $\sigma(\mathrm{Cay}(\mathbb{Z}_{2}^{n},S))\geq\sqrt{|S|}$ for any connection set $S$. In [Potechin-Tsang-20], Potechin and Tsang proved a corresponding result for all Cayley graphs of abelian groups using $|V_{G}|/2$ rather than $\alpha(G)$ as the cutoff for the number of vertices, so they prove that $\sqrt{|S|/2}$ provides a lower bound on sensitivity for bipartite graphs. They conjectured that for every Cayley graph $\mathrm{Cay}(G,S)$, the maximum degree of an induced subgraph on more than half the vertices is at least $\sqrt{|S|/2}$, but Lehner and Verret provided an infinite family of counterexamples [Lehner-Verret-20]. Sensitivity and related properties have also been investigated by several authors for the natural generalization of hypercubes to Hamming graphs $H_{n,q}=\mathrm{Cay}(\mathbb{Z}_{q}^{n},\{\pm e_{1},\dots,\pm e_{n}\})$ for $q\geq 3$ (see Theorem 8.1 in [GarciaMarco-Knauer-22] and [Dong-21, Tandya-22, Potechin-Tsang-24, AFGMK-25]).
@@
The definition of sensitivity of an indexed graph family and three additional infinite families of counterexamples to the Potechin and Tsang conjecture appear in [GarciaMarco-Knauer-22] by Garcìa-Marco and Knauer. All of these families have unbounded degree but sensitivity equal to one, which they refer to as insensitive families. The Hamming graph families $\{H_{n,q}\}_{n=1}^{\infty}$ for $q\geq 3$ and $\{H_{n,q}\}_{q=3}^{\infty}$ are also insensitive since $\Delta(H_{n,q})=(q-1)n$ and for $q\geq 3$ [Tandya-22] constructs an induced subgraph of maximum degree one on $\alpha(H_{n,q})+1$ vertices.
@@
In this paper we formalize definitions from [GarciaMarco-Knauer-22] and [Huang-19], describe graph sensitivity under the join operation and a partition into stable blocks, and apply these results to give general constructions of sensitive and insensitive graph families. We investigate the sensitivity of several explicit, mostly non-regular, families, some of which serve as models for citation, transportation, or neural networks, and find sensitive and insensitive subfamilies occurring within singly- and doubly-indexed families of graphs in a variety of ways. We are aware of two other references that consider sensitivity of nonregular families using a graph operation. Both use a spectral approach as in [Huang-19] to analyze Cartesian products: of connected signed graphs one of which is bipartite in [HLL-21], and of paths in [ZH-24]. More details of the contents are as follows.
@@
In Section 2 after preliminary definitions we present a few examples of sensitive and insensitive graph families, and the definition of $k$-sensitivity, $\sigma_{k}(G)$, of a nonempty graph $G$ for $1\leq k\leq|V_{G}|-\alpha(G)$ as the minimum of the maximum degrees of induced subgraphs on $\alpha(G)+k$ vertices (see the second concluding remark in [Huang-19] and Question (5.1) in [ZH-24]).
@@ 2. Preliminaries
If $G=(V,E)$ is a graph and $S\subseteq V$ is a subset of vertices, the induced subgraph with vertex set $S$ (or subgraph induced by $S$), denoted $G[S]$, is the graph with vertex set $S$, and an edge between $u$ and $v$ if and only if $u$ and $v$ are adjacent in $G$. At times, for simplicity of notation $S$ might be used when $G[S]$ is meant.
@@
The definition of the sensitivity of a graph and what it means for an indexed graph family to be sensitive or insensitive are motivated by Huang’s result [Huang-19] on hypercubes and the discussion in [GarciaMarco-Knauer-22]. To compute $\sigma(G)$ for a nonempty graph it suffices to consider vertex sets with $|S|=\alpha(G)+1$ since the maximum degree can only increase when using a larger vertex set, but all possible such sets must be considered, not just those formed by adding one vertex to a maximum order empty induced subgraph. Defining $\sigma(\overline{K}_{n})$ to be $n$ allows us to state Theorem 3.1 uniformly for all graphs. We do not consider the family of empty graphs to be sensitive or insensitive and in general the condition that $\Delta(G_{n})\to\infty$ as $n\to\infty$ distinguishes insensitive graph families from those where $\sigma(G_{n})$ is bounded because $\Delta(G_{n})$ is bounded.
@@
The initial motivating example of a sensitive family is the hypercube family $Q_{n}$, for which $\sigma(Q_{n})\geq\left\lceil\sqrt{n}\,\right\rceil$ [Huang-19, CFGS-88]. To introduce two simple examples of sensitive non-regular families, recall that the join $G\vee H$ of two graphs $G$ and $H$ is the disjoint union of the two graphs together with all possible edges between them. A natural and useful first example is star graphs.
@@
Taking the cone of the star graphs results in another sensitive family.
@@
$\sigma_{k}(G)\coloneqq\min\{\Delta H\colon H\text{ is an $\alpha(G)+k$-vertex induced subgraph of $G$}\}.$
@@
Note that for all graphs $\sigma_{1}(G)=\sigma(G)$ and for all nonempty graphs we have $\sigma_{k}(G)=\Delta_{\beta}(G)$ for $\beta=\frac{\alpha(G)+k}{|V_{G}|}$ when written in the notation of [GarciaMarco-Knauer-22].
@@
For Hamming graphs with $q\geq 3$ the sensitivity $\sigma(H_{n,q})=1$ (see [Dong-21, Tandya-22, Potechin-Tsang-24]). In [Potechin-Tsang-24] Potechin and Tsang find additional values of $k>1$ such that $\sigma_{k}(H_{n,3})=1$.
@@ 3.1. Join of two graphs
Let $G_{1}$ and $G_{2}$ be graphs with independence numbers $\alpha_{1}\geq\alpha_{2}$ and vertex sets $V_{1}$ and $V_{2}$. Let
@@
with $0\leq j\leq\min\{|V_{2}|,\alpha_{1}+1\}$, whenever they are defined. Then $\sigma(G_{1}\vee G_{2})$ will be the minimum of these $a_{j}$.
@@
For the double graph $G\vee G$ with $G$ nonempty this is immediate from Theorem 3.1 (2) and Lemma 2.12 because $\alpha_{2}=\alpha_{1}$ so the minimum in Theorem 3.1 (2) only involves the two quantities $\sigma(G)$ and $\left\lceil\frac{\alpha(G)+1}{2}\right\rceil$. When $G=\overline{K}_{m}$, then this follows similarly by Theorem 3.1 (1) and Lemma 2.12. In fact, since $\alpha(G\vee G)=\alpha(G)$, repeatedly joining $G$ with itself more than once will yield the same sensitivity as for the double graph of $G$. ∎
@@
If $G$ is a graph and $n\geq 1$, then
$\sigma(G\vee K_{n})=\begin{cases}\sigma(G)\ &\text{if $G$ is nonempty, and}\\ m\ &\text{if $G=\overline{K}_{m}$,}\end{cases}$
and the family $\{G\vee K_{n}\}_{n=1}^{\infty}$ is insensitive while $\{\overline{K}_{m}\vee K_{n}\}_{m=1}^{\infty}$ is sensitive.
@@
If $G$ is nonempty, then using that $\sigma_{k}(K_{n})=k$ for all $n\geq 2$ and $1\leq k\leq n-1$ (Example 2.8) and $K_{n}$ as $G_{2}$ in Theorem 3.1 yields that
@@
Consider the windmill graphs $W_{m,n}=K_{1}\vee mK_{n}$, where $mK_{n}$ is $m$ disjoint copies of $K_{n}$. By Corollary 3.4 we know that when $G$ is nonempty $\sigma(K_{1}\vee G)=\sigma(G)$ and taking the $1$-cone of a nonempty graph preserves sensitivity. Hence, provided $n>1$, all of the subfamilies of the windmill graphs are insensitive with $\Delta(K_{1}\vee mK_{n})=mn$ and $\sigma(K_{1}\vee mK_{n})=\sigma(mK_{n})=1$. The specific windmill graphs $K_{1}\vee mK_{2}$ are often referred to as Dutch windmill or friendship graphs.
@@ 3.2. Joining an empty graph to a nonempty graph
Let $G$ be any graph, let $n\geq 2$, and consider the $n$-cone $\overline{K}_{n}\vee G$. Let
@@
In case (3) since $\alpha(\overline{K}_{n})=|\overline{K}_{n}|\leq\alpha(G)$, instead $\overline{K}_{n}$ serves as $_{2}, GG$ serves as $G_{1}$, and only the first part of Theorem 3.1 applies, which yields the first expression for $\sigma(\overline{K}_{n}\vee G)$. We can also apply Lemma 2.12 to more precisely describe the sensitivity as in the second expression. ∎
@@
We consider both $\overline{K}_{m}\vee G_{n}$ for $G_{n}=P_{n}$, referred to as agave graphs in [Estrada-16], and $G_{n}=C_{n}$, referred to as generalized wheel graphs in [Buckley-Harary-88]. We take $m\geq 3$ and $n\geq 4$ or $n\geq 5$ respectively in order to avoid anomalies in the maximum degree and sensitivity for small graphs. With these assumptions on $m$ and $n$ we have
@@
A complete multipartite graph is regular if and only if each part has the same cardinality since if there were two parts with a different number of vertices, then vertices in those parts would have different degrees. We denote by $K_{n}^{m}$ the complete regular multipartite graph with $m$ parts of order $n$ and find its sensitivity by applying Proposition 3.11 to find $\sigma(K_{n}^{2})=\sigma(\overline{K}_{n}\vee\overline{K}_{n})=\left\lceil\frac{n+1}{2}\right\rceil$. Since also $\Delta(\overline{K}_{n}\vee\cdots\vee\overline{K}_{n})=(m-1)n$, this provides another doubly indexed family with both sensitive and insensitive subfamilies.
@@
The generalized join is also referred to as the joined union [Chudnovsky-24] or Zykov sum [Liao-Aziz-Alaoui-Hou-22]. Many common constructions can be realized as other special cases of generalized joins. For example, $K_{2}[G_{1},G_{2}]=G_{1}\vee G_{2}$ is the standard join, $G[H,\dots,H]=G\circ H$ is the lexicographic product or graph composition of $G$ and $H$, and the complete multipartite graphs in Proposition 3.11 are generalized joins of empty graphs over a complete graph, i.e., of the form $K_{m}[\overline{K}_{n_{1}},\dots,\overline{K}_{n_{m}}]$.
The windmill graphs $W_{m,n}=K_{1}\vee mK_{n}$ in Example 3.6, which can be viewed as the generalized join $S_{m+1}[K_{1},K_{n},\dots,K_{n}]$, were used in [Estrada-16] to explore divergence between often-used transitivity indices and proposed as appropriate models for citation and collaboration networks that exhibit similar divergence properties. In order to extend and expand these results, in [Kooij-19] Kooij considered both what he called type I generalized windmill graphs $K_{m}\vee nH=S_{n+1}[K_{m},H,\dots,H]$ and type II generalized windmill graphs $\overline{K}_{m}\vee nH=S_{n+1}[\overline{K}_{m},H,\dots,H]$, as well as a type III graph that is actually a corona - see Definition 4.17. The $W_{m,n}$ and Types I and II in [Kooij-19] are all subsumed as particular instances of $G[H_{0},H_{1},\dots,H_{n}]$ with $G=S_{n+1}$. The generalized join or joined union $G[H_{1},\dots,H_{n}]$ for any graph $G$ of order $n$ is described as a fourth family of generalized windmill graphs in [Liao-Aziz-Alaoui-Hou-22], where it is noted they more closely relate to real-life transportation networks, and in [Chudnovsky-24] it is pointed out that the graphs $H_{1},\dots,H_{n}$ can be considered to describe the layers of a multilayer network with $G$ controlling how the layers are joined. We consider a few cases of generalized joins for which we can determine sensitivity.
@@
The generalized join of a set of graphs $H_{1},H_{2},\dots,H_{m}$ over an empty graph $\overline{K}_{m}$ is the disjoint union of the $H_{i}$ and hence the sensitivity is
$\sigma(\overline{K}_{m}[H_{1},\dots,H_{m}])=\sigma(H_{1}+\dots+H_{m})=\min\limits_{1\leq i\leq m}\{\sigma(H_{i})\}.$
This generalized join thus corresponds to the partition of a graph into connected components, which is referred to as parallel decomposition (see Section 5)
@@
$\sigma(K_{m}\circ H)=\sigma(H\vee H)=\min\left\{\sigma(H),\left\lceil\frac{\alpha(H)+1}{2}\right\rceil\right\}.$
@@
For the generalized windmill graphs of type I in [Kooij-19], by Corollary 3.4 and Example 3.14 we have
$\sigma(S_{n+1}[K_{m},H_{1},\dots,H_{n}])=\sigma(K_{m}\vee(H_{1}+\dots+H_{n}))=\min\limits_{1\leq i\leq n}\{\sigma(H_{i})\}.$
For certain Type II generalized windmill graphs $S_{n+1}[\overline{K}_{m},H,\dots,H]=\overline{K}_{m}\vee nH$ we can use Theorem 3.7 to more explicitly compute sensitivity because it may be possible to compute $\sigma_{k}(nH)$ by Proposition 2.11. Assume $H$ is not empty because if $H=\overline{K}_{\ell}$ for some $\ell\geq 1$ then $\sigma(\overline{K}_{m}\vee nH)=\sigma(\overline{K}_{m}\vee\overline{K}_{n\ell})$ is known by Corollary 3.2. Note that in general
@@
Sensitive families can thus be constructed by either taking a sequence of $n$-cones of any graph, as in Corollary 3.8, or by fixing $m$ and taking the $m$-cone of a family of sensitive graphs, as in Corollary 3.17. For instance, $\{\overline{K}_{m}\vee Q_{2n+1}^{+}\}_{n=1}^{\infty}$, $\{\overline{K}_{m}\vee K_{1,n}\}_{n=1}^{\infty}$, and $\{\overline{K}_{m}\vee K_{1,1,n}\}_{n=1}^{\infty}$ are all sensitive. Insensitive families can be constructed by joining any graph with the sequence of complete graphs, as in Corollary 3.4, or by building certain types of generalized joins - e.g., a sequence of Type II generalized windmill graphs with an increasing number of pendant graphs, or in which one of the pendant graphs has sensitivity $1$ as in Example 3.16. Other operations, like taking the $1$-cone of a nonempty graph or doubling any graph, preserve sensitivity or insensitivity.

### v3

@@ 1. Introduction
The idea of graph sensitivity traces back to work related to Boolean functions and the hypercube graph ([13, 6]), but has been formalized and studied for other graphs following Huang’s clever and concise resolution in [17] of the sensitivity conjecture for Boolean functions [25]. The hypercube graph, $Q_{n}$, which has vertices the $n$-tuples in $\mathbb{Z}_{2}^{n}$ with two vertices adjacent if they differ in exactly one coordinate, is bipartite so its independence number is $\alpha(Q_{n})=\frac{1}{2}2^{n}=2^{n-1}$. Huang proved the smallest maximum degree of an induced subgraph of $Q_{n}$ on $\alpha(Q_{n})+1$ vertices is at least $\sqrt{n}$ and the construction in [6] implies it is equal to $\left\lceil\sqrt{n}\,\right\rceil$. Huang further suggests studying, for other highly symmetric graphs, the smallest maximum degree of an induced subgraph of a graph $G$ on $\alpha(G)+1$ vertices, defined as the sensitivity $\sigma(G)$ in [11] by Garcìa-Marco and Knauer.
@@
Since $Q_{n}$ is the Cayley graph on $\mathbb{Z}_{2}^{n}$ with connection set $S=\{e_{1},\dots,e_{n}\}$ it is natural to investigate the sensitivity of other Cayley graphs. For example, in [1] Alon and Zheng proved $\sigma(\mathrm{Cay}(\mathbb{Z}_{2}^{n},S))\geq\sqrt{|S|}$ for any connection set $S$. In [26], Potechin and Tsang proved a corresponding result for all Cayley graphs of abelian groups using $|V_{G}|/2$ rather than $\alpha(G)$ as the cutoff for the number of vertices, so they prove that $\sqrt{|S|/2}$ provides a lower bound on sensitivity for bipartite graphs. They conjectured that for every Cayley graph $\mathrm{Cay}(G,S)$, the maximum degree of an induced subgraph on more than half the vertices is at least $\sqrt{|S|/2}$, but Lehner and Verret provided an infinite family of counterexamples in [23]. Sensitivity and related properties have also been investigated by several authors for the natural generalization of hypercubes to Hamming graphs $H_{n,q}=\mathrm{Cay}(\mathbb{Z}_{q}^{n},\{\pm e_{1},\dots,\pm e_{n}\})$ for $q\geq 3$ (see Theorem 8.1 in [11] as well as [8, 29, 27, 2]).
@@
The definition of sensitivity of an indexed graph family and three additional infinite families of counterexamples to the Potechin and Tsang conjecture appear in [11] by Garcìa-Marco and Knauer. All of these families have unbounded degree but sensitivity equal to one, which they refer to as insensitive families. The Hamming graph families $\{H_{n,q}\}_{n=1}^{\infty}$ for $q\geq 3$ and $\{H_{n,q}\}_{q=3}^{\infty}$ are also insensitive since $\Delta(H_{n,q})=(q-1)n$ and for $q\geq 3$ Tandya constructs an induced subgraph of maximum degree one on $\alpha(H_{n,q})+1$ vertices in [29].
@@
In this paper we formalize definitions from [11] and [17], describe graph sensitivity under the join operation and a partition into stable blocks, and apply these results to give general constructions of sensitive and insensitive graph families. We investigate the sensitivity of several explicit, mostly non-regular, families, some of which serve as models for citation, transportation, or neural networks, and find sensitive and insensitive subfamilies occurring within singly- and doubly-indexed families of graphs in a variety of ways. We are aware of two other references that consider sensitivity of nonregular families using a graph operation. Both use a spectral approach as in [17] to analyze Cartesian products: of connected signed graphs one of which is bipartite in [16], and of paths in [30].
@@
More details of the contents of this paper are as follows. In Section 2 after preliminary definitions we present a few examples of sensitive and insensitive graph families, and the definition of $k$-sensitivity, $\sigma_{k}(G)$, of a nonempty graph $G$ for $1\leq k\leq|V_{G}|-\alpha(G)$ as the minimum of the maximum degrees of induced subgraphs on $\alpha(G)+k$ vertices (see the second concluding remark in [17] and Question (5.1) in [30]).
@@ 2. Preliminaries
If $G=(V,E)$ is a graph and $S\subseteq V$ is a subset of vertices, the induced subgraph with vertex set $S$ (or subgraph induced by $S$), denoted $G[S]$, is the graph with vertex set $S$, and an edge between $u$ and $v$ if and only if $u$ and $v$ are adjacent in $G$. At times, for simplicity of notation $S$ might be used if the context makes it clear that $G[S]$ is meant.
@@
The definition of the sensitivity of a graph and what it means for an indexed graph family to be sensitive or insensitive are motivated by Huang’s result [17] on hypercubes and the discussion in [11]. To compute $\sigma(G)$ for a nonempty graph it suffices to consider vertex sets with $|S|=\alpha(G)+1$ since the maximum degree can only increase when using a larger vertex set, but all possible such sets must be considered, not just those formed by adding one vertex to a maximum order empty induced subgraph. Defining $\sigma(\overline{K}_{n})$ to be $n$ allows us to state Theorem 3.1 uniformly for all graphs. However, we do not consider the family of empty graphs to be sensitive or insensitive and in general the condition that $\Delta(G_{n})\to\infty$ as $n\to\infty$ distinguishes insensitive graph families from those where $\sigma(G_{n})$ is bounded because $\Delta(G_{n})$ is bounded.
@@
The initial motivating example of a sensitive family is the hypercube family $Q_{n}$, for which $\sigma(Q_{n})\geq\left\lceil\sqrt{n}\,\right\rceil$ [17, 6]. To introduce two simple examples of sensitive non-regular families, recall that the join $G\vee H$ of two graphs $G$ and $H$ is the disjoint union of the two graphs together with all possible edges between them. A natural and useful first example is star graphs.
@@
Taking the cone of the star graphs results in another sensitive family with sensitivity $n$.
@@
$\sigma_{k}(G)\coloneqq\min\{\Delta H\colon H\text{ is an $(\alpha(G)+k)$-vertex induced subgraph of $G$}\}.$
@@
Note that for all graphs $\sigma_{1}(G)=\sigma(G)$ and for all nonempty graphs we have $\sigma_{k}(G)=\Delta_{\beta}(G)$ for $\beta=\frac{\alpha(G)+k}{|V_{G}|}$ when written in the notation of [11].
@@
For Hamming graphs with $q\geq 3$ the sensitivity $\sigma(H_{n,q})=1$ (see [8, 29, 27]). In [27] Potechin and Tsang find additional values of $k>1$ such that $\sigma_{k}(H_{n,3})=1$.
@@ 3.1. Join of two graphs
Let $G_{1}$ and $G_{2}$ be graphs with independence numbers $\alpha_{1}\geq\alpha_{2}$ and vertex sets $V_{1}$ and $V_{2}$. Define
@@
with $0\leq j\leq\min\{|V_{2}|,\alpha_{1}+1\}$ whenever they are defined. Then $\sigma(G_{1}\vee G_{2})$ will be the minimum of these $a_{j}$.
@@
For the double graph $G\vee G$ with $G$ nonempty this follows from Theorem 3.1 (2) and Lemma 2.12 because $\alpha_{2}=\alpha_{1}=\alpha(G)$ so the minimum in Theorem 3.1 (2) only involves the two quantities $\sigma(G)$ and $\left\lceil\frac{\alpha(G)+1}{2}\right\rceil$. When $G=\overline{K}_{m}$, then this follows by Corollary 3.2. In fact, since $\alpha(G\vee G)=\alpha(G)$, repeatedly joining $G$ with itself more than once will yield the same sensitivity as for the double graph of $G$. ∎
@@
If $G$ is a graph and $n\geq 1$, then $\sigma(G\vee K_{n})=\sigma(G)$ and the family $\{G\vee K_{n}\}_{n=1}^{\infty}$ is insensitive while $\{\overline{K}_{m}\vee K_{n}\}_{m=1}^{\infty}$ is sensitive.
@@
If $G$ is nonempty, then using that $\sigma_{k}(K_{n})=k$ for all $n\geq 2$ and $1\leq k\leq n-1$ by Example 2.8 and $K_{n}$ as $G_{2}$ in Theorem 3.1 yields that
@@
Consider the windmill graphs $W_{m,n}=K_{1}\vee mK_{n}$, where $mK_{n}$ is $m$ disjoint copies of $K_{n}$. By Corollary 3.4 we know that $\sigma(K_{1}\vee G)=\sigma(G)$ and taking the $1$-cone of a graph preserves sensitivity. Hence, provided $n>1$, all of the subfamilies of the windmill graphs are insensitive with $\Delta(K_{1}\vee mK_{n})=mn$ and $\sigma(K_{1}\vee mK_{n})=\sigma(mK_{n})=1$ while $\sigma(K_{1}\vee mK_{1})=\sigma(K_{1}\vee\overline{K}_{m})=m$ as in Example 2.4. The specific windmill graphs $K_{1}\vee mK_{2}$ are often referred to as Dutch windmill or friendship graphs.
@@ 3.2. Joining an empty graph to a nonempty graph
Let $G$ be any graph and consider the $n$-cone $\overline{K}_{n}\vee G$ for $n\geq 1$. Let
@@
In case (3) since $\alpha(\overline{K}_{n})=|\overline{K}_{n}|\leq\alpha(G)$, instead $\overline{K}_{n}$ serves as $_{2}, GG$ serves as $G_{1}$, and only the first part of Theorem 3.1 applies, which yields the first expression for $\sigma(\overline{K}_{n}\vee G)$. We can also apply Lemma 2.12 to more precisely describe the sensitivity as in the second expression. In particular, if $n=1$ then we recover $\sigma(K_{1}\vee G)=\sigma(G)$. ∎
@@
We consider both $\overline{K}_{m}\vee G_{n}$ for $G_{n}=P_{n}$, referred to as agave graphs in [9], and $G_{n}=C_{n}$, referred to as generalized wheel graphs in [4]. We take $m\geq 3$ and $n\geq 4$ or $n\geq 5$ respectively in order to avoid anomalies in the maximum degree and sensitivity for small graphs. With these assumptions on $m$ and $n$ we have
@@
A complete multipartite graph is regular if and only if each part has the same cardinality since if there were two parts with a different number of vertices, then vertices in those parts would have different degrees. We denote by $K_{n}^{m}$ the complete regular multipartite graph with $m$ parts of order $n$ and find its sensitivity by applying Proposition 3.11 to find $\sigma(K_{n}^{m})=\sigma(K_{n}^{2})=\sigma(\overline{K}_{n}\vee\overline{K}_{n})=\left\lceil\frac{n+1}{2}\right\rceil$. Since also $\Delta(K_{n}^{m})=\Delta(\overline{K}_{n}\vee\cdots\vee\overline{K}_{n})=(m-1)n$, this provides another doubly indexed family with both sensitive and insensitive subfamilies.
@@
The generalized join is also referred to as the joined union [5] or Zykov sum [24]. Many common constructions can be realized as other special cases of generalized joins. For example, $K_{2}[G_{1},G_{2}]=G_{1}\vee G_{2}$ is the standard join, $G[H,\dots,H]=G\circ H$ is the lexicographic product or graph composition of $G$ and $H$, and the complete multipartite graphs in Proposition 3.11 are generalized joins of empty graphs over a complete graph, i.e., of the form $K_{m}[\overline{K}_{n_{1}},\dots,\overline{K}_{n_{m}}]$.
The windmill graphs $W_{m,n}=K_{1}\vee mK_{n}$ in Example 3.6, which can be viewed as the generalized join $S_{m+1}[K_{1},K_{n},\dots,K_{n}]$, were used in [9] to explore divergence between often-used transitivity indices and proposed as appropriate models for citation and collaboration networks that exhibit similar divergence properties. In order to extend and expand these results, in [18] Kooij considered both what he called type I generalized windmill graphs $K_{m}\vee nH=S_{n+1}[K_{m},H,\dots,H]$ and type II generalized windmill graphs $\overline{K}_{m}\vee nH=S_{n+1}[\overline{K}_{m},H,\dots,H]$, as well as a type III graph that is actually a corona - see Definition 4.17. The $W_{m,n}$ and Types I and II in [18] are all subsumed as particular instances of $G[H_{0},H_{1},\dots,H_{n}]$ with $G=S_{n+1}$. The generalized join or joined union $G[H_{1},\dots,H_{n}]$ for any graph $G$ of order $n$ is described as a fourth family of generalized windmill graphs in [24], where it is noted they more closely relate to real-life transportation networks, and in [5] it is pointed out that the graphs $H_{1},\dots,H_{n}$ can be considered to describe the layers of a multilayer network with $G$ controlling how the layers are joined. We consider a few cases of generalized joins for which we can determine sensitivity.
@@
The generalized join of a set of graphs $H_{1},H_{2},\dots,H_{m}$ over an empty graph $\overline{K}_{m}$ is the disjoint union of the $H_{i}$. The sensitivity is thus
[…]
because adding an isolated vertex to a nonempty graph does not change the sensitivity. When all of the $H_{i}$ are nonempty this generalized join corresponds to the partition of a graph into connected components, which is referred to as the parallel decomposition (see Section 5).
@@
$\sigma(K_{m}\circ H)=\sigma(H\vee H\vee\cdots\vee H)=\min\left\{\sigma(H),\left\lceil\frac{\alpha(H)+1}{2}\right\rceil\right\}.$
@@
For the generalized windmill graphs of type I in [18], by Corollary 3.4 and Example 3.14 we have $S_{n+1}[K_{m},H_{1},\dots,H_{n}]=K_{m}\vee(H_{1}+\dots+H_{n})$ and hence
[…]
For certain Type II generalized windmill graphs $S_{n+1}[\overline{K}_{m},H,\dots,H]=\overline{K}_{m}\vee nH$ we can use Theorem 3.7 to more explicitly compute sensitivity when it is possible to compute $\sigma_{k}(nH)$ using Proposition 2.11. Assume $H$ is not empty because if $H=\overline{K}_{\ell}$ for some $\ell\geq 1$ then $\sigma(\overline{K}_{m}\vee nH)=\sigma(\overline{K}_{m}\vee\overline{K}_{n\ell})$ is known by Corollary 3.2. Note that in general
@@
Sensitive families can thus be constructed by either taking a sequence of $n$-cones of any graph, as in Corollary 3.8, or by fixing $m$ and taking the $m$-cone of a family of sensitive graphs, as in Corollary 3.17. For instance, $\{\overline{K}_{m}\vee Q_{2n+1}^{+}\}_{n=1}^{\infty}$, $\{\overline{K}_{m}\vee K_{1,n}\}_{n=1}^{\infty}$, and $\{\overline{K}_{m}\vee K_{1,1,n}\}_{n=1}^{\infty}$ are all sensitive. Insensitive families can be constructed by joining any graph with the sequence of complete graphs, as in Corollary 3.4, or by building certain types of generalized joins - e.g., a sequence of Type II generalized windmill graphs with an increasing number of pendant graphs, or in which one of the pendant graphs has sensitivity $1$ as in Example 3.16. Other operations, like taking the $1$-cone of a graph or doubling any graph, preserve sensitivity or insensitivity.